OpenStudy (anonymous):

plz help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! how many solutions does the equation have? 4(2x+3)=2(3x-4) A. one solution B. infinite solutions C. no solution

OpenStudy (yummy123):

hold on

OpenStudy (yummy123):

A is correct

OpenStudy (anonymous):

thx so much i fan you and i will medal you too

OpenStudy (yummy123):

thanks

OpenStudy (anonymous):

np

OpenStudy (yummy123):

do you want me to explain

OpenStudy (anonymous):

yeah it might help me understand better

OpenStudy (yummy123):

the question is 4(2x+3) = 2(3x-4) 8x+12 = 6x-8 2x = -20 x=-10 which is leads to A

OpenStudy (yummy123):

does that help

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

What is the best way to classify each equation? Column A 22-3x+7x=4(x+5) A. identity contradiction neither

OpenStudy (welshfella):

simplify both sides and see what you get Note: an identity is true for all values of x

OpenStudy (anonymous):

6(2x-3)=3(3x-5)

OpenStudy (welshfella):

a contradiction does not make sense

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

im not good at this double sided equation stuff

OpenStudy (welshfella):

22 - 3x + 7x = 22 + 4x 4(x + 5) = ?

OpenStudy (welshfella):

multiply x + 5 by 4 = 4*x + 4*5 can you continue?

OpenStudy (welshfella):

4 8 x is simply 4x

OpenStudy (anonymous):

4*x+20

OpenStudy (welshfella):

4 * x

OpenStudy (anonymous):

so 4+x

OpenStudy (anonymous):

imean 4=x

OpenStudy (welshfella):

ok so we have 4x + 22 = 4x + 20 identity, contradiction or neither?

OpenStudy (welshfella):

can that be true?

OpenStudy (anonymous):

yes?

OpenStudy (welshfella):

lets say x = 1 we then have 4*1 + 22 = 4*1 + 20 so 26 = 24

OpenStudy (welshfella):

26=24!!????

OpenStudy (anonymous):

so its neither contradiction or identity

OpenStudy (welshfella):

wrong its a contradiction - it doesn't make sense. There can be no solution to it.

OpenStudy (anonymous):

ok thx

OpenStudy (welshfella):

26 is obviusly not equal to 24!!

OpenStudy (anonymous):

yeah im just stupid when it come to double sided equations

OpenStudy (welshfella):

you need to practice them .

OpenStudy (welshfella):

6(2x-3)=3(3x-5) 6*2x - 6*3 = 3*3x + 3*-5 12x - 18 = 9x - 15 this one looks like an equation with one solution First subtract 9x from each side to get all x's on the left 12x - 18 = 9x - 15 -9x -9x 3x - 18 = -15 now add 18 to both sides can you try that?